OpenStudy (anonymous):

jeff bikes to work in 3/4 hrs. his bus ride takes 1/4 hr. the bus goes 20 mph faster than the bikes. how far is work????

OpenStudy (amistre64):

too far apparently; jeff needs a job that can afford him a car

OpenStudy (anonymous):

way to help

OpenStudy (amistre64):

since the distance never changes; we can equate both parts to d

OpenStudy (amistre64):

d = speed * time

OpenStudy (amistre64):

d = Bs(3/4) d = (Bs+20)(1/4)

OpenStudy (anonymous):

(Speed_bike+20)*15=Speed_bike*45

OpenStudy (anonymous):

Speed bike =10mph Speed bus = 30mph

OpenStudy (amistre64):

Bs 3/4 = Bs 1/4 + 5 Bs 2/4 = 5 Bs 2 = 20 Bs = 10 mph

OpenStudy (anonymous):

Distance = 7.5 m

OpenStudy (amistre64):

10*3/4 = 30/4 = 7' 1/2 miles

OpenStudy (anonymous):

Solve the following for bike, the bikes speed.\[\text{bike} \frac{3}{4}=(\text{bike}+20)\frac{1}{4}\]bike = 10 3/4*10 = 7.5 miles